A rock falls from rest a vertical distance of 1.32 meter to the surface of a planet in 0.63 second. The magnitude of the acceleration due to gravity on the planet is approximately

1
Expert's answer
2022-01-24T13:15:54-0500

Given:

"h=1.32\:\rm m"

"t=0.63\: \rm s"

"g-?"


The time of free falling

"t=\sqrt{2h/g}"

Hence, the acceleration due to the gravity

"g=\frac{2h}{t^2}=\rm\frac{2*1.32\: m}{(0.63\: s)^2}=6.65\: m/s^2"